OpenStudy (flvskidd):

HELP! Ahmad's gas tank is 310 full. After he buys 7 gallons of gas, it is 45 full. How many gallons can Ahmad's tank hold?

OpenStudy (flvskidd):

@agent0smith

OpenStudy (flvskidd):

@PartyShotz2409

OpenStudy (flvskidd):

@foreverjb

OpenStudy (flvskidd):

@ProdigyNation55

OpenStudy (flvskidd):

@phi

OpenStudy (flvskidd):

@radar

OpenStudy (agent0smith):

Screenshot.

OpenStudy (muscrat1233):

can you attach a screenshot of the question?

OpenStudy (flvskidd):

yes sry

OpenStudy (muscrat1233):

lol its fine

OpenStudy (flvskidd):

the question changed but its the same idea

OpenStudy (muscrat1233):

lol okay

OpenStudy (muscrat1233):

okay so what you want to do is make it so that the two fractions have the same denominator

OpenStudy (muscrat1233):

so what you can do is make 1/2 into 5/10

OpenStudy (muscrat1233):

so the container is 8/10 full, or 80% full

OpenStudy (flvskidd):

I made it 5/10 how can i tell how many gallons

OpenStudy (flvskidd):

@muscrat1233

OpenStudy (muscrat1233):

im working it hang on

OpenStudy (peachpi):

If x is tank volume in gallons \[\frac{3}{10}x+5=\frac{1}{2}x\]

OpenStudy (flvskidd):

sry i dont get it

OpenStudy (flvskidd):

thx though

OpenStudy (agent0smith):

It went from 3/10 full, to 5/10 full (since 1/2 is 5/10) Which means they added 2/10 of water. They also told you that the 2/10 ths they added was 5 gallons. So if x is the total tank size, then\[\large \frac{ 2 }{ 10 } x = 5\]

OpenStudy (phi):

Are you still stuck?